solve the no-linear equation:

x+y =7
x^2+ y^2=25

Answer by checkley71(8403)

x=7-y substiruting 7-y for x in the second equation we get
(7-y)~2+y~2=25 or 49-7y-7y+y~2+y~2=25 or 49-14y+y~2+y~2=25 or 2y~2-14y+24=0 or
dividing all terms by 2 we get x~2-7x-12=0 or (y-4)(y-3)=0 thus y=4 & y=3